Rosnan Industries' 2017 and 2016 balance sheets and income statements are shown below. All of the balance of Cash and Equivalents is an operating asset.

Balance Sheets:
2017 2016
Cash and equivalents $100   $85  
Accounts receivable 275   200  
Inventories 375   250  
      Total current assets $750   $635  
Net plant and equipment 2,000   1,490  
Total assets $2,750   $2,125  
Accounts payable $150   $85  
Accruals 75   50  
Notes payable 150   75  
      Total current liabilities $375   $210  
Long-term debt 450   290  
Common stock 1,225   1,225  
Retained earnings 700   400  
Total liabilities and equity $2,750   $2,125  


Income Statements:
2017 2016
Sales $2,000   $1,500  
Operating costs excluding depreciation 1,250   1,000  
EBITDA $750   $500  
Depreciation and amortization 100   75  
EBIT $650   $425  
Interest 62   45  
EBT $588   $380  
Taxes (40%) 235   152  
Net income $353   $228  
Dividends paid $53   $48  
Addition to retained earnings $300   $180  
Shares outstanding 190   190  
Price $20.83   $18.33  
WACC 12.00%     


Using the financial statements above, what is Rosnan's 2017 market value added (MVA)? Round your answer to the nearest dollar. Do not round intermediate calculations.
$

Using the financial statements given earlier, what is Rosnan's 2017 economic value added (EVA)? Round your answer to the nearest cent. Do not round intermediate calculations.
$
Answer #1

MVA = Market Value of Firm - Capital invested in firm = (Market value of Stock + Market value of Debt) - (Book Value of Stock + Book value of Debt ) = (190*20.833 +450) -( 1225 + 450) = 2733.27 or 2733


EVA = NOPAT - Capital Charge = EAT + Interest*(1-Tax) - WACC * Capital Employed = 353+ 62*(1-40%) - 12% * 2750 = 60.2

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well